# Muon g-2 and lepton flavor violation in supersymmetric GUTs

### v1

@@ 1 Introduction
In addition to providing a solution to the hierarchy problem by shielding the Higgs boson mass from quadratically divergent radiative corrections, supersymmetry (SUSY) is a well motivated extension of the Standard Model (SM) since it also offers a plausible dark matter (DM) candidate. In addition, the supersymmetric partners of the SM particles provide new contributions to the muon anomalous magnetic moment (muon $g-2$) to better fit with the current experimental measurements Muong-2:2023cdq ; Muong-2:2021ojo ; Muong-2:2006rrc . Finally, a successful unification of the SM gauge couplings in its minimal supersymmetric extension (MSSM) motivates one to build SUSY grand unified theories (GUTs) that are spontaneously broken at the grand unification scale ($M_{{\rm GUT}}\simeq 2\times 10^{16}$ GeV). A solution to the muon $g-2$ problem together with a compelling explanation of the Higgs boson mass and plausible DM candidates can lead to mass relations among the SUSY particles which can be tested in the current experiments. For instance, while the muon $g-2$ favors a relatively light SUSY mass spectrum, the observed Higgs boson mass of 125 GeV ATLAS:2012yve ; CMS:2013btf requires a fairly heavy set of SUSY particles. Some GUT relations among the soft gaugino masses can alleviate this challenge by predicting heavy gluinos, which enhance the squark masses through the renormalization group equations (RGEs), while the sleptons remain relatively light Chakraborti:2021dli ; Baer:2021aax ; Aboubrahim:2021xfi ; Wang:2021bcx ; Han:2020exx ; Altin:2017sxx ; Li:2021pnt ; Ellis:2021zmg ; Athron:2021iuf ; Chakraborti:2021bmv ; Endo:2021zal ; Iwamoto:2021aaf ; Baum:2021qzx ; Frank:2021nkq ; Heinemeyer:2021opc ; Akula:2013ioa ; Gomez:2022qrb . These schemes can be compatible with flavor-mixing scalar soft terms at the GUT scale, where the scalar masses are mostly induced by flavor-independent large gaugino masses through the RGEs Gomez:2010ga . In these cases, flavor violation is suppressed in the quark sector, while the lepton sector can still yield interesting phenomenology, especially since it is favored by the muon $g-2$ solution.
@@
The SUSY contributions to muon $g-2$ can arise at one-loop from the electroweak gauginos, Higgsinos and sleptons through the diagrams shown in Fig. 1 Martin:2001st ; Giudice:2012pf ; Moroi:1995yh , and these contributions potentially fill the gap between the latest measurements Muong-2:2023cdq and the muon $g-2$ predctions SM Aoyama:2020ynm . However, the current experimental constraints, especially from the Higgs boson mass, can exclude the required relatively large contributions in the simplest SUSY models such as CMSSM. In addition, when leptons and sleptons mass matrices are not aligned in the same superfield basis, Lepton Flavor Violating (LFV) processes can occur through very similar topologies as shown in Fig. 2. In this section, we consider the muon $g-2$ and LFV processes and emphasize the interplay between them.
@@ 2.1 SUSY contribution to muon g-2
We will work within the $4-2-2$ framework presented in Ref. Gomez:2022qrb . The relevant SUSY parameters at the GUT scale are listed in Table 1. As previously shown, in this framework the muon $g-2$ solutions favor heavy gluinos, while the LSP can still be a good DM candidate with masses lighter than about 500 GeV. The heavy gluinos enhance the squark masses through RGEs, thus leading to a significant splitting between the squarks and slepton masses. This can easily be seen by using the approximate RGE solutions provided in Refs. Ibanez:1984vq ; Martin:1997ns that show an approximate relation between the GUT scale parameters and the low energy mass terms.
$m_{\tilde{Q}_{i}}^{2}\!\!\!$ $=$ $\!\!\!m_{L}^{2}+K_{3}+K_{2}+{\frac{1}{36}1}K_{1},$
@@
Although this approximation is only valid for cases where first and second generation Yukawa couplings can be neglected still it provides a good picture of the mass splitting among sqarks and sfermions. Furthermore, these equations can show how a possible flavor dependence of the scalar soft terms can be transmitted to the low energy observables. For instance, we observe that the sfermion masses at the TeV scale have a flavor independent component arising from gauginos in addition to the contribution from the soft scalar masses, that can be flavor dependent. This may induce misalignments between the fermion Yukawa and the sfermion mass matrices. Typically, in SUSY models the flavor mixture of the soft terms results in fermionic flavor violating decays, very constrained by the strong bounds on flavor changing neutral currents (FCNC). However, in the $4-2-2$ framework we find that the large gluino masses suitable for explaining muon $g-2$, and the lower masses for $M_{1},M_{2}$ that can explain DM, can suppress such processes on the quarks while being less restrictive on the leptonic sector. For instance, according to Eqs. 4 only diagonal mass terms grow with gaugino masses, thus the flavour mixing in the squark mass matrices are washed out by the large gluino contribution to the diagonal ones, such that their contribution to processes like $b\rightarrow s\gamma$ Olive:2008vv is not significant. On the other hand, the moderate values we find for the lighter gauginos do not suppress the effect of the slepton mass mixing terms, leading to the prediction of charged lepton violating (cLFV) decays on the range of the experimental searches Gomez:2010ga .
@@
In the current work, we use gauge group $4-2-2$ without imposing Left-Right (LR) symmetry (for a detailed description see Gomez:2020gav ; Gomez:2022qrb and references therein). Such symmetries can emerge from $SO(10)$ breaking by the Higgs fields in $210_{H}$ representation Babu:2016bmy . Assuming type I see-saw, cLFV will be induced below the GUT scale. The running of the slepton masses from $M_{GUT}$ to $M_{R}$ is affected by the Dirac neutrino Yukawa coupling matrix $Y_{\nu}$, that can be of the same order as the fermion Yukawa couplings. Therefore, flavor-changing terms for the slepton masses are introduced since $Y_{\nu}$ cannot be diagonalised in the same basis as the charged lepton Yukawa matrix $Y_{l}$. For instance, in a basis where $Y_{l}$ is diagonal at the leading-log approximation Hisano:1995cp , the non-diagonal terms of the lepton mass matrix take the following form at scale $M_{R}$:
@@
$m_{f}^{2}=\begin{pmatrix}1&\epsilon^{q_{L}^{12}}&\epsilon^{q_{L}^{13}}\\ \epsilon^{q_{L}^{12}}&1&\epsilon^{q_{L}^{23}}\\ \epsilon^{q_{L}^{12}}&\epsilon^{q_{L}^{23}}&1\\ \end{pmatrix}\times m_{L}^{2},\hskip 14.22636ptm_{f^{c}}^{2}=\begin{pmatrix}1&\epsilon^{q_{R}^{12}}&\epsilon^{q_{R}^{13}}\\ \epsilon^{q_{R}^{12}}&1&\epsilon^{q_{R}^{23}}\\ \epsilon^{q_{R}^{12}}&\epsilon^{q_{R}^{23}}&1\\ \end{pmatrix}\times m_{R}^{2}~{},$ (10)
where $m_{L}$ and $m_{R}$ are the soft sypersymmetry breaking (SSB) masses for the left- and right-handed sfermions. These matrices are induced both in the squark the slepton sectors. However, because of the heavy squarks driven by the heavy gluinos through RGEs, the flavor mixing in these matrices is suppressed and their flavor mixing in essentially coincides with the CKM matrix. On the other hand, even though the charged lepton Yukawa matrix remains diagonal in this basis, the non-diagonal form of the slepton mass-squared matrix can induce the cLFV processes such as $l_{i}\rightarrow l_{j}\gamma$ at the loop level, as shown diagrammatically in Fig. 2.
@@ 3 Fundamental Parameters and Experimental Constraints
We perform random scans in the fundamental parameter space of 4-2-2 with broken LR symmetry by using SPheno-4.0.4 Porod:2003um ; Porod:2011nf , which is implemented by SARAH-4.14.4 Staub:2008uz . The values for the input parameters at the GUT level used in the 4-2-2 model described above are:
@@
Here, $m_{L}$ is the SSB mass term for the left-handed sfermions and $M_{2L}$, $M_{3}$ stand for the SSB masses of $SU(2)_{L}$ and $SU(3)_{c}$ gauginos, respectively. The LR symmetry breaking in the sfermions and gauginos is controlled by ${\rm x}_{{\rm LR}}$ and ${\rm y}_{{\rm LR}}$, namely $m_{R}={\rm x}_{{\rm LR}}m_{L}$ and $M_{2R}=y_{LR}M_{2L}$, where $M_{2R}$ denotes the SSB mass of $SU(2)_{R}$ gaugino. Similarly, $x_{d}$ and $x_{u}$ parametrize the non-universality in the SSB masses at $M_{{\rm GUT}}$ of the MSSM Higgs fields, $m_{H_{d}}^{2}=x_{d}m_{L}^{2}$ and $m_{H_{u}}^{2}=x_{u}m_{L}^{2}$. In addition, we vary the universal trilinear scalar interaction term $A_{0}$ by requiring that the magnitude of its ratio to $m_{L}$ is not greater than 3, to avoid the color/charge breaking minima of the scalar potential Ellwanger:1999bv ; Camargo-Molina:2013qva ; Camargo-Molina:2013sta . In scanning the parameter space we use the Metropolis-Hastings algorithm Belanger:2009ti ; Baer:2008jn , and generate the solutions by following flat priors Trotta:2008bp .
As indicated in the previous section, we introduce LFV in three different scenarios. One is simplified seesaw Type 1 where the flavor mixing among left sleptons is induced by RGEs. The other two scenarios mix the slepton flavors at scales above $M_{{\rm GUT}}$. The texture is chosen so that, for $\epsilon=0.05$, it provides values for BR($\mu\rightarrow e\gamma$) comparable to the seesaw case, while for $\epsilon=0.2$ it may strongly restrict the GUT values of the soft masses, allowing only models comparable with ”no-scale” SUSY Lahanas:1986uc ; Ellis:2013nka ; Li:2016bww ; Ford:2019kzv ; Li:2021cte . We impose the the bounds ParticleDataGroup:2022pth :
@@
In our search, we keep the points that satisfy the constraints:
$\begin{array}[]{c}m_{h}=123-127~{}{\rm GeV}~{},\\ m_{\tilde{g}}\geq 2.1~{}{\rm TeV}~{}(800~{}{\rm GeV}~{}{\rm if~{}it~{}is~{}NLSP})~{},\\ 0.8\times 10^{-9}\leq{\rm BR}(B_{s}\rightarrow\mu^{+}\mu^{-})\leq 6.2\times 10^{-9}\;(2\sigma)~{},\\ 2.99\times 10^{-4}\leq{\rm BR}(B\rightarrow X_{s}\gamma)\leq 3.87\times 10^{-4}\;(2\sigma)~{},\\ 0.114\leq\Omega_{{\rm CDM}}h^{2}\leq 0.126~{}.\end{array}$ (16)
In addition to requiring the solutions to be consistent with these constraints, we also select those which predict muon $g-2$ within the $2\sigma$ band of its experimentally measured values and also include the interplay between the cLFV processes and muon $g-2$ contributions. Finally, we only accept solutions which are compatible with the LSP neutralino relic density either consistent or lower than the Planck measurements Planck:2015fie . The DM constraints together with the muon $g-2$ condition imposes a severe restriction on the possible DM scenarios, that are limited to the following cases:
  • Refer to caption

### v3

@@ 1 Introduction
In addition to providing a solution to the hierarchy problem by shielding the Higgs boson mass from quadratically divergent radiative corrections, supersymmetry (SUSY) is a well motivated extension of the Standard Model (SM) since it also offers a plausible dark matter (DM) candidate. In addition, the supersymmetric partners of the SM particles provide new contributions to the muon anomalous magnetic moment (muon $g-2$) to better fit with the current experimental measurements Muong-2:2023cdq ; Muong-2:2024hpx ; Muong-2:2021ojo ; Muong-2:2006rrc . Finally, a successful unification of the SM gauge couplings in its minimal supersymmetric extension (MSSM) motivates one to build SUSY grand unified theories (GUTs) that are spontaneously broken at the grand unification scale ($M_{{\rm GUT}}\simeq 2\times 10^{16}$ GeV). A solution to the muon $g-2$ problem together with a compelling explanation of the Higgs boson mass and plausible DM candidates can lead to mass relations among the SUSY particles which can be tested in the current experiments. For instance, while the muon $g-2$ favors a relatively light SUSY mass spectrum, the observed Higgs boson mass of 125 GeV ATLAS:2012yve ; CMS:2013btf requires a fairly heavy set of SUSY particles. Some GUT relations among the soft gaugino masses can alleviate this challenge by predicting heavy gluinos, which enhance the squark masses through the renormalization group equations (RGEs), while the sleptons remain relatively light Chakraborti:2021dli ; Baer:2021aax ; Aboubrahim:2021xfi ; Wang:2021bcx ; Han:2020exx ; Altin:2017sxx ; Li:2021pnt ; Ellis:2021zmg ; Athron:2021iuf ; Chakraborti:2021bmv ; Endo:2021zal ; Iwamoto:2021aaf ; Baum:2021qzx ; Frank:2021nkq ; Heinemeyer:2021opc ; Akula:2013ioa ; Gomez:2022qrb ; Ellis:2024ijt . These schemes can be compatible with flavor-mixing scalar soft terms at the GUT scale, where the scalar masses are mostly induced by flavor-independent large gaugino masses through the RGEs Gomez:2010ga . In these cases, flavor violation is suppressed in the quark sector, while the lepton sector can still yield interesting phenomenology, especially since it is favored by the muon $g-2$ solution.
@@
The SUSY contributions to muon $g-2$ can arise at one-loop from the electroweak gauginos, Higgsinos and sleptons through the diagrams shown in Fig. 1 Martin:2001st ; Giudice:2012pf ; Moroi:1995yh , and these contributions potentially fill the gap between the latest measurements Muong-2:2024hpx and the muon $g-2$ predctions SM Aoyama:2020ynm . However, the current experimental constraints, especially from the Higgs boson mass, can exclude the required relatively large contributions in the simplest SUSY models such as CMSSM. In Gomez:2022qrb we have shown that MSSM models derived from $4-2-2$ scheme described above can explain muon $g-2$. Here, we will discuss in addition that when leptons and sleptons mass matrices are not aligned in the same superfield basis, Lepton Flavor Violating (LFV) processes can occur through very similar topologies as shown in Fig. 2. In this section, we consider the muon $g-2$ and LFV processes and emphasize the interplay between them.
2.1 A ${4-2-2}$ SUSY-GUT framework.
In this subsection we review some aspects of the $4-2-2$ symmetry relevant for extracting the MSSM framework that we will use in our computations. The $4-2-2$ is one of the symmetries containing the SM Pati:1974yy ; Lazarides:1980tg ; Kibble:1982ae that can be preserved in different breaking schemes of $SO(10)$ Babu:1992ia ; Anderson:1993fe ; Drees:1986vd ; Kawamura:1993uf ; Kolda:1995iw ; Miller:2012vn ; Babu:2005ui ; Milagre:2024wcg , for instance, through the vacuum expectation values (VEVs) of the Higgs fields residing either in $54_{H}$ or $210_{H}$ Higgs multiplets. The breaking through the VEV in $54_{H}$ preserves the so called $C-$parity Kibble:1982dd ; Lazarides:1985my , which transforms the left-handed and right-handed fields into each other. However, if one implements the $SO(10)$ breaking through the $210_{H}$ VEV, the $C-$parity is broken Babu:2016bmy . In this work we will adopt the latter scheme, assuming left and right fields belong to disconnected $4-2-2$ representations.
The breaking of $4-2-2$ to the MSSM can be driven at the GUT scale by the Higgs fieldsKing:1997ia :
$\bar{H}$ $=$ $(\bar{4},1,\bar{2})=(u^{c}_{H},d^{c}_{H},\nu^{c}_{H},e^{c}_{H})$
$H$ $=$ $(4,1,2)=(\bar{u}^{c}_{H},\bar{d}^{c}_{H},\bar{\nu}^{c}_{H},\bar{e}^{c}_{H})$ (1)
[…]
$Y=\sqrt{\dfrac{3}{5}}I_{3R}+\sqrt{\dfrac{2}{5}}(B-L)~{},$ (2)
here $I_{3R}$ and $B-L$ represent the diagonal generators of $SU(2)_{R}$ and $SU(4)_{c}$, respectively. Taking into account supersymmetry breaking yields the following relation among the three soft-susy breaking (SSB) gaugino masses at the GUT scale:
$M_{1}=\dfrac{3}{5}M_{2R}+\dfrac{2}{5}M_{4}~{},$ (3)
where $M_{2R}$ and $M_{4}$ denote the gaugino mass terms for $SU(2)_{R}$ and $SU(4)_{c}$, respectively, and $M_{3}=M_{4}$ at $M_{{\rm GUT}}$. If the LR breaking in the gaugino sector is parametrised as $M_{2R}=y_{LR}M_{2L}$, where $M_{2L}$ is the SSB mass of $SU(2)_{L}$ gaugino, the SSB gaugino masses become independent. This is crucial in order to find a supersymmetric solution to the muon $g-2$, since typically models with universal gaugino masses add a small contribution to it. This is a consequence of the large SUSY spectrum resulting from the gluino mass bound. In contrast, non-universal gaugino masses contribute selectively to a SUSY spectrum where large squark masses coexists with lighter sleptons that can explain the muon $g-2$ anomaly.
The squarks and leptons reside in the representations:
$F_{i}=(4,2,1)_{i}$ $=$ $\left(\begin{array}[]{cccc}u_{r}&u_{g}&u_{b}&\nu_{l}\\ d_{r}&d_{g}&d_{b}&e_{l}\end{array}\right)_{i},$ (6)
$\bar{F}_{i}=(\bar{4},2,1)_{i}$ $=$ $\left(\begin{array}[]{cccc}u^{c}_{r}&u^{c}_{g}&u^{c}_{b}&\nu^{c}_{l}\\ d^{c}_{r}&d^{c}_{g}&d^{c}_{b}&e^{c}_{l}\end{array}\right)_{i},$ (9)
where $i=1,2,3$ is the family index and $r,g,b,l$ denote the 4 colors of $SU(4)_{C}$.
The broken LR symmetry, in general, implies non-universal masses for the soft left- and right-handed matter scalar masses which can be quantified as $m_{R}\equiv x_{LR}\cdot m_{L}$, where $m_{R}$ ($m_{L}$) denotes the SSB mass of the right-handed (left-handed) fields. In this work, we assume that the $4-2-2$ symmetry is broken to the MSSM gauge group near the GUT scale, and thus we require the solutions to maintain approximately unification of the MSSM gauge couplings at $M_{{\rm GUT}}$. The $4-2-2$ particle spectrum contains right-handed neutrinos, which play a role in explaining neutrino oscillations established by the current experiments Super-Kamiokande:2010orq . However, an extenision of the MSSM with tiny neutrino masses by a see-saw mechanism requires either very heavy right-handed neutrinos or negligible couplings between the MSSM particles and the right-handed neutrinos ($Y_{\nu}\lesssim 10^{-7}$) Coriano:2014wxa ; Khalil:2010zza ; Abbas:2007ag , and so the right-handed neutrinos effectively decouple from the MSSM spectrum. Nevertheless, they can leave signals in LFV processes as we will see later.
Reagarding the MSSM Higgs sector, both Higgs doublets, $H_{u}$, $H_{d}$ can be contained in the $4-2-2$ representation $h=(1,2,\bar{2})$, allowing for Yukawa coupling unification Gomez:2020gav . However, in this work we consider a more general scenario where the MSSM Higgs fields are realized as superpositions of the appropriate Higgs fields residing in different representations mentioned in the $SO(10)$ and $4-2-2$ breakings. Hence, we relax the condition of Yukawa unification and keep free the ratio of the Higgs vev’s, $tan\beta=v_{u}/v_{d}$. Also, we assume independent SSB mass terms for the MSSM Higgs fields.
Thus, the model below $M_{{\rm GUT}}$ turns out to be MSSM with some relations among the SSB parameters at the GUT scale:
  • Gaugino masses. Although there is a $4-2-2$ realtion among gaugino masses, the identification $M_{3}=M_{4}$ and $M_{2}=M_{2L}$, together with breaking of the $L/R$ symmetry that implies $M_{2R}=y_{LR}\cdot M_{2L}$, results in independent MSSM gaugino masses.
  • Scalar masses. Common masses for squarks and leptons but distinction between the left and right fields $m_{R}\equiv x_{LR}\cdot m_{L}$. Different Higgs mass parameters: $m_{H_{u}}=x_{u}\cdot M_{L}$, $m_{H_{d}}=x_{d}\cdot M_{L}$.
To summarize, the $4-2-2$ GUT symmetry determines a MSSM with the following free parameters:
$M_{4},M_{2L},y_{LR},M_{L},x_{LR},x_{u},x_{d},A_{0}/M_{L},tan\beta,$ (10)
where the masses can go up to the TeV scale while the parameters $y_{LR}$, $x_{LR}$, $x_{u}$, $x_{d}$ and $A_{0}/M_{L}$ are of order unity. Using the relations explained above, it is possible to identify the extracted MSSM framework with
$M_{3},M_{2},M_{1},M_{L},M_{R},m_{H_{u}},m_{H_{d}},A_{0},tan\beta,$ (11)
as free parameters.
The model is complemented with a mechanism that can introduce LFV without interfering with the flavor conserving SUSY contributions to muon $g-2$ and DM studied in Gomez:2022qrb . This can be done by assuming a see-saw mechanism that generates LFV slepton masses below the GUT scale, or additional U(1) family symmetries that explain the flavor structure of the Yukawa couplings and are broken at a scale above the $M_{GUT}$ introducing family dependent soft sfermion masses that still preserve the $4-2-2$ symmetry. In the resulting MSSM, the fermions and sfermions mass matrices are not aligned in the same superfield basis leading to a SUSY contribution to flavor violating processes in charged leptons Calibbi:2017uvl ; Vicente:2015cka and quarks, like B and kaon systems Isidori:2006qy ; Altmannshofer:2009ne ; Crivellin:2017gks ; Endo:2017ums , imposing strong bounds on the flavor structure of the models Olive:2008vv ; Gomez:1998wj . However, the RG induced gauino contribution to soft masses is flavor independent such that potential flavor mixing soft masses introduced at GUT can be substantially reduced at the EW scale Gomez:2010ga . Furthermore, in the $4-2-2$ framework under consideration, despite the fact that squarks and sleptons get common SSB terms, the splitting of the gaugino masses induce squark masses much larger than the sleptons ones. Therefore, in this framework we can find significant SUSY contributions to LFV without enhancing FV in the quark sector.
@@ 2.2 SUSY contribution to muon g-2
We will work within the $4-2-2$ framework presented in Ref. Gomez:2022qrb . The relevant SUSY parameters at the GUT scale are listed in Table 1. As previously shown, in this framework the muon $g-2$ solutions favor heavy gluinos, while the LSP can still be a good DM candidate with masses lighter than about 500 GeV. The heavy gluinos enhance the squark masses through RGEs, thus leading to a significant splitting between the squarks and slepton masses. This can easily be seen by using the approximate RGE solutions provided in Refs. Ibanez:1984vq ; Martin:1997ns that show an approximate relation between the GUT scale parameters and the low energy mass terms. Although the results presented in subsequent sections are obtained by running the full RGE set by using the public code SPheno Porod:2003um ; Porod:2011nf , these approximations allow us to understand the relations amoung SUSY masses in the different models and thus, the relevance of the SUSY contributions to the processes under consideration.
$m_{\tilde{Q}_{i}}^{2}\!\!\!$ $=$ $\!\!\!m_{L}^{2}+K_{3}+K_{2}+{\frac{1}{36}}K_{1},$
@@
Although this approximation is only valid for cases where first and second generation Yukawa couplings can be neglected still it provides a good picture of the mass splitting among sqarks and sfermions. Furthermore, these equations can show how a possible flavor dependence of the scalar soft terms can be transmitted to the low energy observables. For instance, we observe that the sfermion masses at the TeV scale have a flavor independent component arising from gauginos in addition to the contribution from the soft scalar masses, that can be flavor dependent. This may induce misalignments between the fermion Yukawa and the sfermion mass matrices.
Typically, in SUSY models the flavor mixture of the soft terms results in fermionic flavor violating decays, heavily constrained by the strong experimental bounds Gomez:1998wj ; King:2004tx ; Olive:2008vv . However, in the $4-2-2$ framework we find that the large gluino masses suitable for explaining muon $g-2$, and the lower masses for $M_{1},M_{2}$ that can explain DM, helps suppress such processes in the quark sector while being less restrictive in the leptonic sector. For instance, although all the sfermions may get FV masses of the same order at the GUT scale, according to Eqs. 15 only the diagonal mass terms grow with gaugino masses, and thus the flavour mixing in the squark mass matrices are washed out by the large gluino contribution to the diagonal ones. Therefore, the SUSY contribution to sensitive systems like kaon decays Aebischer:2022vky is not significant. On the other hand, the moderate values we find for the lighter gauginos do not suppress the effect of the slepton mass mixing terms, leading to the prediction of charged lepton violating (cLFV) decays on the range of the experimental searches Gomez:2010ga .
@@
In the current work, we use gauge group $4-2-2$ without imposing Left-Right (LR) symmetry (for a detailed description see Gomez:2020gav ; Gomez:2022qrb and references therein). Such symmetries can emerge from $SO(10)$ breaking by the Higgs fields in $210_{H}$ representation Babu:2016bmy . Assuming I see-saw, cLFV will be induced below the GUT scale. The running of the slepton masses from $M_{GUT}$ to $M_{R}$ is affected by the Dirac neutrino Yukawa coupling matrix $Y_{\nu}$, that can be of the same order as the fermion Yukawa couplings. Therefore, flavor-changing terms for the slepton masses are introduced since $Y_{\nu}$ cannot be diagonalised in the same basis as the charged lepton Yukawa matrix $Y_{l}$. For instance, in a basis where $Y_{l}$ is diagonal at the leading-log approximation Hisano:1995cp , the non-diagonal terms of the lepton mass matrix take the following form at scale $M_{R}$:
@@
$m_{f}^{2}=\begin{pmatrix}1&\epsilon^{q_{L}^{12}}&\epsilon^{q_{L}^{13}}\\ \epsilon^{q_{L}^{12}}&1&\epsilon^{q_{L}^{23}}\\ \epsilon^{q_{L}^{12}}&\epsilon^{q_{L}^{23}}&1\\ \end{pmatrix}\times m_{L}^{2},\hskip 14.22636ptm_{f^{c}}^{2}=\begin{pmatrix}1&\epsilon^{q_{R}^{12}}&\epsilon^{q_{R}^{13}}\\ \epsilon^{q_{R}^{12}}&1&\epsilon^{q_{R}^{23}}\\ \epsilon^{q_{R}^{12}}&\epsilon^{q_{R}^{23}}&1\\ \end{pmatrix}\times m_{R}^{2},$ (21)
where $m_{L}$ and $m_{R}$ are the soft sypersymmetry breaking (SSB) masses for the left- and right-handed sfermions. These matrices are induced both in the squark the slepton sectors. However, because of the heavy squarks driven by the heavy gluinos through RGEs, the flavor mixing in these matrices is suppressed and their flavor mixing in essentially coincides with the CKM matrix. Therefore, sensitive processes like rare kaon decays Aebischer:2022vky are not significantly affected in this models. On the other hand, even though the charged lepton Yukawa matrix remains diagonal in this basis, the non-diagonal form of the slepton mass-squared matrix can induce the cLFV processes such as $l_{i}\rightarrow l_{j}\gamma$ at the loop level, as shown diagrammatically in Fig. 2.
@@ 3 Fundamental Parameters and Experimental Constraints
In Gomez:2022qrb we have shown the existence of models with soft terms derived from the $4-2-2$ model that can simultaneously explain muon $g-2$ without running into conflict with the observed DM relic density. The input parameters at the GUT level for these models correspond to values of the $4-2-2$ parameters in eq. 11 as follows:
@@
The low energy spectrum and phenomenological constraints are obtained using SPheno-4.0.4 Porod:2003um ; Porod:2011nf , which is implemented by SARAH-4.14.4 Staub:2008uz . This is done by using the MSSM parameters listed in eq. 10 and using MSSM and the see-saw type I packages included in the SPheno distribution. The selected models are compatible with EWSB and satisfy the GUT unification condition as described in Gomez:2022qrb . In addition, we vary the universal trilinear scalar interaction term $A_{0}$ by requiring that the magnitude of its ratio to $m_{L}$ is not greater than 3, to avoid the color/charge breaking minima of the scalar potential Ellwanger:1999bv ; Camargo-Molina:2013qva ; Camargo-Molina:2013sta . In scanning the parameter space we use the Metropolis-Hastings algorithm Belanger:2009ti ; Baer:2008jn , and generate the solutions by following flat priors Trotta:2008bp .
In our search, we keep the points that satisfy the constraints:
$\begin{array}[]{c}m_{h}=123-127~{}{\rm GeV}~{},\\ m_{\tilde{g}}\geq 2.1~{}{\rm TeV}~{}(800~{}{\rm GeV}~{}{\rm if~{}it~{}is~{}NLSP})~{},\\ 0.8\times 10^{-9}\leq{\rm BR}(B_{s}\rightarrow\mu^{+}\mu^{-})\leq 6.2\times 10^{-9}\;(2\sigma)~{},\\ 2.99\times 10^{-4}\leq{\rm BR}(B\rightarrow X_{s}\gamma)\leq 3.87\times 10^{-4}\;(2\sigma)~{}.\end{array}$ (24)
The Standard Model (SM) Higgs boson mass is computed by SPheno which runs three-loop RGEs by considering the effective SM Higgs potential between $M_{Z}$ and $M_{SUSY}$ and imposing the two-scale matching condition at $M_{SUSY}$ Staub:2017jnp . The SPheno outputs are transferred to micrOMEGAs Belanger:2020gnr , which calculates the relic density and the scattering cross-sections of the DM candidates. We consider the 5-$\sigma$ Planck range Planck:2015fie :
$0.114\leq\Omega_{{\rm CDM}}h^{2}\leq 0.126~{}.$ (25)
In general, we keep models that are not in contradiction with these bounds. That is, models that predict DM relic density below the upper bound in the above equation. However, we will present some of our results for the subsets of models predicting DM predictions within the Planck range.
In contrast to the study presented in Gomez:2022qrb , here we perform a more selective scan accepting only models that explain muon $g-2$ at $2\sigma$. The discrepancy between the combined BNL and Fermilab results Muong-2:2024hpx and the data-driven theoretical value is Aoyama:2020ynm :
$\Delta a_{\mu}=(24.9\pm 4.8)\times 10^{-10}.$ (26)
The impact on the SUSY models from more restrictive ranges, for instance like the ones arising from lattice calculations, are studied in detail in Ref.Ellis:2024ijt .
As indicated in the previous section, we introduce LFV in three different scenarios. One is a simplified seesaw type I, where the flavor mixing among left sleptons is induced by the RGEs. The other two scenarios mix the slepton flavors at scales above $M_{{\rm GUT}}$. The texture is chosen so that, for $\epsilon=0.05$, it provides values for BR($\mu\rightarrow e\gamma$) comparable to the seesaw case, while for $\epsilon=0.2$ it may strongly restrict the GUT values of the soft masses, allowing only models comparable with ”no-scale” SUSY Lahanas:1986uc ; Ellis:2013nka ; Li:2016bww ; Ford:2019kzv ; Li:2021cte ; Ellis:2020mno . We impose the the bounds ParticleDataGroup:2022pth :
@@
The DM constraints together with the muon $g-2$ condition and the cLFV bounds impose a severe restriction on the possible DM scenarios, that are limited to the following cases:
  • Refer to caption Refer to caption Refer to caption